OpenStudy (anonymous):

A tomato fits in a 10- in.-long golden rectangle. What is the tomato's width to the nearest tenth of an inch?

OpenStudy (anonymous):

length/width = phi length/phi = width

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

the golden ratio is approximately 1.6180339887.... which is often represented by phi so for the rectangle to be a golden ratio it holds that length divided by width must be equal to phi so your length divided by phi is your width. or L/W=phi so L=Phi*W and W=L/Phi